Determine the extension of the spring when it has a spring constant k= 60 N/m with 25 J of potential energy.


1
Expert's answer
2021-03-31T16:14:32-0400
"PE=\dfrac{1}{2}kx^2,""x=\sqrt{\dfrac{2PE}{k}}=\sqrt{\dfrac{2\cdot25\ J}{60\ \dfrac{N}{m}}}=0.91\ m."